# CRPS-Based Targeted Sequential Design with Application in Chemical Space

### v1

@@ Abstract
Sequential design of real and computer experiments via Gaussian Process (GP) models has proven useful for parsimonious, goal-oriented data acquisition purposes. In this work, we focus on acquisition strategies for a GP model that needs to be accurate within a predefined range of the response of interest. Such an approach is useful in various fields including synthetic chemistry, where finding molecules with particular properties is essential for developing useful materials and effective medications. GP modeling and sequential design of experiments have been successfully applied to a plethora of domains, including molecule research. Our main contribution here is to use the threshold-weighted Continuous Ranked Probability Score (CRPS) as a basic building block for acquisition functions employed within sequential design. We study pointwise and integral criteria relying on two different weighting measures and benchmark them against competitors, demonstrating improved performance with respect to considered goals. The resulting acquisition strategies are applicable to a wide range of fields and pave the way to further developing sequential design relying on scoring rules.
@@ 1 Introduction
Scoring rules (Gneiting and Raftery,, 2007) are widely used tools for assessing probabilistic forecasts by comparing a predictive distribution with a single value (the materialized response). Scoring rules evaluate both the statistical consistency between predictions and observations (calibration) and the sharpness of the prediction. An interesting category of scoring rules for goal-oriented prediction evaluation consists of weighted scoring rules, allowing to emphasize particular outcomes by incorporating a weighting measure (Gneiting and Ranjan,, 2011). In this study, we focus on the Continuous Ranked Probability Score (CRPS; Matheson and Winkler, 1976), for which Allen et al., 2023b outline three potential approaches to use a weighting measure. Here, we consider the threshold-weighted CRPS introduced by Matheson and Winkler, (1976) and Gneiting and Ranjan, (2011), which has been widely used for evaluating forecasts of extreme weather events (e.g., Lerch and Thorarinsdottir,, 2013). To reflect the special importance of predictions in/near the excursion set $\Gamma$, we consider both an indicator weight on $\Gamma$, and a Gaussian weight centered at $t$.
@@
We showcase and compare the derived criteria and resulting sequential strategies on chemical test cases revolving around the Photoswitch dataset (Thawani et al.,, 2020). GP models have demonstrated indeed promising results in predicting molecular properties, as highlighted for instance in the work of Thawani et al., (2020), Deringer et al., (2021) and Griffiths et al., (2023). Thereby, the choice of kernel heavily depends on the molecule representation used to encode chemical structures. We consider molecules represented by molecular fingerprints (Christie et al.,, 1993; Johnson and Maggiora,, 1990), which encode the presence or absence of characteristic substructures using binary digits. For fingerprint representations, the Tanimoto kernel (Gower,, 1971) is widely utilized (e.g., Ralaivola et al.,, 2005; Miyao and Funatsu,, 2019; Tripp et al.,, 2023), favored due to its simplicity, efficiency, and effectiveness in measuring the overlap between the presence or absence of molecular features. By combining the Tanimoto kernel with our targeted CRPS-based criteria, we demonstrate how scoring rules can be leveraged for efficient sequential design towards uncovering molecules of interest.
The structure of the remaining paper is as follows: Section 2 reviews the fundamentals of GP modeling, excursion set estimation and targeted sequential design. While Section 3 is dedicated to the threshold-weighted CRPS and the development of the acquisition criteria based thereon, the numerical tests are introduced in Section 4 and the results are presented in Section 5. Section 6 wraps up the main body with a discussion and concluding words.
@@ 2.1 Gaussian process modeling and excursion set estimation
[…]
@@ 3.1 Threshold-weighted CRPS
Weighted scoring rules enable a targeted assessment of forecasts. The threshold-weighted CRPS (Matheson and Winkler,, 1976; Gneiting and Ranjan,, 2011) extends the CRPS by considering a weighting measure $\nu$, which emphasizes the values of interest:
$\mathrm{CRPS}_{\nu}(F,y)=\int_{-\infty}^{\infty}[F(u)-\mathbbm{1}\{y\leq u\}]^{2}\,\mathrm{d}\nu(u).$ (7)
While $\nu$ may potentially be taken as any Borel measure on $\mathbb{R}$, in our excursion setting, we consider two types of weighting measures (characterized by their densities with respect to the Lebesgue measure $\lambda$). The first weighting measure $\nu_{1}$ only considers values above the threshold $t$ and is characterized by $\mathrm{d}\nu_{1}(u)=\mathbbm{1}\{t\leq u\}\mathrm{d}\lambda(u)$. The second weighting measure $\nu_{2}$ admits a Gaussian probability density centered on the threshold $t$ and with variance $\sigma_{\nu}^{2}>0$: $\mathrm{d}\nu_{2}(u)=\frac{1}{\sigma_{\nu}}\phi\left(\frac{u-t}{\sigma_{\nu}}\right)\mathrm{d}\lambda(u)$, where $\phi:\mathbb{R}\rightarrow(0,\infty),\phi(v):=\frac{1}{\sqrt{2\pi}}\exp^{-\frac{v^{2}}{2}}$.
@@
For a predictive distribution $\Phi_{\mu,\sigma^{2}}$, it holds with weighting measure $\nu_{1}$ that
$\mathrm{CRPS}_{\nu_{1}}(\Phi_{\mu,\sigma^{2}},y)=\sigma\left[-\widetilde{t}\Phi(\widetilde{t})^{2}+\widetilde{y^{t}}\left(2\Phi(\widetilde{y^{t}})-1\right)+\left(2\phi\left(\widetilde{y^{t}}\right)-2\phi\left(\widetilde{t}\right)\Phi(\widetilde{t})\right)-\frac{1}{\sqrt{\pi}}\left(1-\Phi(\widetilde{t}\sqrt{2})\right)\right],$
@@
For the weighting measure $\nu_{2}$ (with mean $t$ and variance $\sigma_{\nu}^{2}$), one obtains
$\mathrm{CRPS}_{\nu_{2}}(\Phi_{\mu,\sigma^{2}},y)$ $=\Phi\left(\begin{pmatrix}0\\ 0\end{pmatrix};\begin{pmatrix}\mu-t\\ \mu-t\end{pmatrix},\begin{pmatrix}\sigma_{\nu}^{2}+\sigma^{2}&\sigma_{\nu}^{2}\\ \sigma_{\nu}^{2}&\sigma_{\nu}^{2}+\sigma^{2}\end{pmatrix}\right)$
$\quad\quad-2\Phi\left(\begin{pmatrix}0\\ 0\end{pmatrix};\begin{pmatrix}\mu-t\\ y-t\end{pmatrix},\begin{pmatrix}\sigma_{\nu}^{2}+\sigma^{2}&\sigma_{\nu}^{2}\\ \sigma_{\nu}^{2}&\sigma_{\nu}^{2}\end{pmatrix}\right)+\Phi\left(\frac{t-y}{\sigma_{\nu}}\right),$
@@
See Appendix A.3.1. ∎
For illustration purposes, Figure 1 depicts the discussed versions of the CRPS for a Gaussian $\Phi_{\mu,\sigma^{2}}$ (black solid line), showing the area under the integral for the CRPS (a and d), $\mathrm{CRPS}_{\nu_{1}}$ (b and e), and $\mathrm{CRPS}_{\nu_{2}}$ (c and f) across two scenarios: In the first row (a-c) we consider $y<t<\mu$, referring to a point which does not lie in the excursion set ($y<t$), but is falsely predicted to be in the excursion set ($\mu>t$, see Eq. 2). In the second row (d-f), we use $\mu<t<y$, such that the point does lie in the excursion set ($y>t$), but is falsely predicted to be not ($\mu<t$). Notably, both the (unweighted) CRPS and $\mathrm{CRPS}_{\nu_{2}}$ are symmetric, yielding the same value for both scenarios. Contrarily, the $\mathrm{CRPS}_{\nu_{1}}$ has a higher value in the second scenario.
Figure 1: Illustrations of the $\mathrm{CRPS}\left(\Phi_{\mu,\sigma^{2}},y\right)$ and $\mathrm{CRPS}_{\nu}\left(\Phi_{\mu,\sigma^{2}},y\right)$ for a Gaussian CDF $\Phi_{\mu,\sigma^{2}}(u)$ (black solid line), an observation $y$ (solid indicator line) and a threshold $t$. In the first row (a-c) we have $y=-3,t=0,\mu=3$ and in the second row (d-f) we have $\mu=-3,t=0,y=3$; we use $\sigma=2$ for the predictive Gaussian. The individual plots display the area under the integral for the CRPS (a and d), $\mathrm{CRPS}_{\nu_{1}}$ (b and e) and $\mathrm{CRPS}_{\nu_{2}}$ (c and f; $\sigma_{\nu}=1.5$), with the dashed lines indicating the weighting measures $\nu_{1}$ and $\nu_{2}$. The numbers refer to the value and expected value of the respective criterion.
@@ 3.2 Pointwise selection criteria
Given that the threshold-weighted CRPS evaluates both accuracy and uncertainty of the GP model with respect to the excursion set of interest, it is a promising basis for an acquisition criterion. For a pointwise criterion, we consider the threshold-weighted CRPS of all the candidate points, and add the one to the training set which has the highest score (Eq. 3). However, in its original form for evaluation (Eq. 7), the threshold-weighted CRPS relies on the materialized value $y$ in the candidate point, which is unavailable during data acquisition. Therefore, we consider an expected version, where the expectation is taken with respect to the current predictive distribution. This is also known as the expected score or entropy function for proper scoring rules (Gneiting and Raftery,, 2007). The expected threshold-weighted CRPS with weighting measure $\nu$ for predictive CDF $F$ is defined as,
$\mathrm{CRPS}_{\nu}(F)=\int_{-\infty}^{\infty}\mathrm{CRPS}_{\nu}(F,y^{\prime})\mathrm{d}F(y^{\prime}).$ (8)
The following theorem provides an alternative representation of this expected threshold-weighted CRPS and derives explicit formulas for Gaussian predictive distributions with weighting measures $\nu_{1}$ and $\nu_{2}$.
@@
For a predictive CDF $F(\cdot)$ and a weighting measure $\nu$, it holds,
$\mathrm{CRPS}_{\nu}(F)=\int F(u)\left(1-F(u)\right)\mathrm{d}\nu(u).$
Particularly, for $F=\Phi_{\mu,\sigma^{2}}$ and $\mathrm{d}\nu_{1}(u)=\mathbbm{1}\{t\leq u\}\mathrm{d}\lambda(u)$, one obtains,
$\mathrm{CRPS}_{\nu_{1}}(\Phi_{\mu,\sigma^{2}})=\sigma\Big{[}\widetilde{t}\Phi(\widetilde{t})^{2}-\widetilde{t}\Phi(\widetilde{t})+\frac{1}{\sqrt{\pi}}-\frac{1}{\sqrt{\pi}}\Phi(\widetilde{t}\sqrt{2})+2\phi(\widetilde{t})\Phi(\widetilde{t})-\phi(\widetilde{t})\Big{]},$
@@
$\mathrm{CRPS}_{\nu_{1}}(\Phi_{\mu,\sigma^{2}})=\sigma\mathbb{E}\left[\left(\widetilde{N}-\max\left(N,\frac{t-\mu}{\sigma}\right)\right)_{+}\right],$
for $N,\widetilde{N}\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$. For $\mathrm{d}\nu_{2}(u)=\frac{1}{\sigma_{\nu}}\phi\left(\frac{u-t}{\sigma_{\nu}}\right)\mathrm{d}\lambda(u)$, we get,
$\mathrm{CRPS}_{\nu_{2}}(\Phi_{\mu,\sigma^{2}})$ $=\Phi\left(\begin{pmatrix}0\\ 0\end{pmatrix};\begin{pmatrix}\mu-t\\ t-\mu\end{pmatrix},\begin{pmatrix}\sigma_{\nu}^{2}+\sigma^{2}&-\sigma_{\nu}^{2}\\ -\sigma_{\nu}^{2}&\sigma_{\nu}^{2}+\sigma^{2}\end{pmatrix}\right).$
@@
Appendix A.3.2; the first result generalizes the corresponding expression for the traditional CRPS (Gneiting and Raftery,, 2007). ∎
To use this for pointwise selection criteria (Eq. 3), we can set $G_{n}(\boldsymbol{x})=\mathrm{CRPS}_{\nu}(\Phi_{\mu,\sigma^{2}})$, where $\mu=m_{n}(\boldsymbol{x})$ and $\sigma^{2}=k_{n}(\boldsymbol{x,\boldsymbol{x}})$. To better understand the resulting behavior, we return to Figure 1, in which we also denote the expected score values. Both CRPS and $\mathrm{CRPS}_{\nu_{2}}$ maintain symmetry in their expected scores (a, c, d, f), while $\mathrm{CRPS}_{\nu_{1}}$ remains asymmetric but with a reversed orientation. In the original scoring rule relying on $y$, a missed point within the excursion set results in a higher score. In the expected score, a point mistakenly classified as being within the excursion set yields a higher expected score. In general, the $\mathrm{CRPS}_{\nu_{1}}$ is higher for cases where $\Phi_{\mu,\sigma^{2}}$ has $m>t$, regardless of whether the prediction is correct (since the correct outcome is unknown). Figure 2a shows this relationship for $\Phi_{\mu,\sigma^{2}}$ with a fixed $t=0$. Both the values of $\mathrm{CRPS}_{\nu_{1}}$ and $\mathrm{CRPS}_{\nu_{2}}$ increase with higher uncertainty ($\sigma$) for the same $\mu$. However, while $\mathrm{CRPS}_{\nu_{1}}$ continuously increases with rising $\mu$ when $\sigma$ is fixed, $\mathrm{CRPS}_{\nu_{2}}$ remains symmetric around the threshold $t$ for varying $\mu$.
Figure 2: Illustrations of the expected $\mathrm{CRPS}_{\nu_{1}}$ and $\mathrm{CRPS}_{\nu_{2}}$ for a Gaussian $\Phi_{\mu,\sigma^{2}}$ depending on the predictive mean $\mu$: (a) Expected $\mathrm{CRPS}_{\nu_{1}}$ and $\mathrm{CRPS}_{\nu_{2}}$ ($\sigma_{\nu}=1.5$) and (b) Expected $\mathrm{CRPS}_{\nu_{2}}$ for different values of $\sigma_{\nu}$ ($\sigma=2$). In both plots, the vertical solid line indicate the threshold $t=0$.
While the indicator density of $\mathrm{CRPS}_{\nu_{1}}$ is fixed (although a shift would be thinkable), the density of $\mathrm{CRPS}_{\nu_{2}}$ is influenced by the parameter $\sigma_{\nu}$, which controls the width of the weight around the threshold $t$. Figure 2b illustrates the dependence of $\mathrm{CRPS}_{\nu_{2}}$ on the predictive mean $\mu$ for different values of $\sigma_{\nu}$ with $t=0$. A lower value of $\sigma_{\nu}$ places greater emphasis on the region near the threshold, while a higher value results in more uniform values across different $\mu$.

### v2

@@ Abstract
Sequential design of real and computer experiments via Gaussian Process (GP) models has proven useful for parsimonious, goal-oriented data acquisition purposes. In this work, we focus on acquisition strategies for a GP model that needs to be accurate within a predefined range of the response of interest. Such an approach is useful in various fields including synthetic chemistry, where finding molecules with particular properties is essential for developing useful materials and effective medications. GP modeling and sequential design of experiments have been successfully applied to a plethora of domains, including molecule research. Our main contribution here is to use the threshold-weighted Continuous Ranked Probability Score (CRPS) as a basic building block for acquisition functions employed within sequential design. We study pointwise and integral criteria relying on two different weighting measures and benchmark them against competitors, demonstrating improved performance with respect to considered goals. The resulting acquisition strategies are applicable to a wide range of fields and pave the way to further developing sequential design relying on scoring rules. Supplemental material to augment this article is available online.
@@ 1 Introduction
Scoring rules (Gneiting and Raftery,, 2007) are widely used tools for assessing probabilistic forecasts by comparing a predictive distribution with a single value (the materialized response). Scoring rules evaluate both the statistical consistency between predictions and observations (calibration) and the sharpness of the prediction. A scoring rule is called proper if no alternative forecast distribution can yield a lower expected score than the true data-generating distribution. It is strictly proper if the expected score is uniquely minimized by the true data-generating distribution. Because proper scoring rules promote careful evaluations and are firmly rooted in fundamental decision-theoretic principles, they are regarded as essential tools in scientific practice (Gneiting and Raftery,, 2007; Bröcker and Smith,, 2007; Thorarinsdottir et al.,, 2013). Furthermore, links between scoring rules and divergences have been actively investigated in the literature of forecast evaluation (Thorarinsdottir et al.,, 2013). In this study, we focus on the widely used Continuous Ranked Probability Score (CRPS; Matheson and Winkler, 1976). There are several advantages that motivate using CRPS (Gneiting et al.,, 2007; Waghmare and Ziegel,, 2025): Echoing with the above, the CRPS has been proven to define a proper scoring rule, and to be strictly proper relative to probability measures that have finite first moment. Furthermore, the CRPS is distance sensitive, in the sense that it rewards forecasts that allocate probability mass closer to the realized observation; and it enjoys theoretical uniqueness as it is the only scoring role satisfying translation invariance, homogeneity, and kernel representability.
An interesting category of scoring rules for goal-oriented prediction evaluation consists of weighted scoring rules, allowing to emphasize particular outcomes by incorporating a weighting measure (Gneiting and Ranjan,, 2011). For the CRPS, Allen et al., 2023b outline three potential approaches to use a weighting measure. Here, we consider the threshold-weighted CRPS introduced by Matheson and Winkler, (1976) and Gneiting and Ranjan, (2011), which has been widely used for evaluating forecasts of extreme weather events (e.g., Lerch and Thorarinsdottir,, 2013). To reflect the special importance of predictions in/near the excursion set $\Gamma$, we consider both an indicator weight on $\Gamma$, and a Gaussian weight centered at $t$.
@@
We showcase and compare the derived criteria and resulting sequential strategies on chemical test cases revolving around the Photoswitch dataset (Thawani et al.,, 2020). In their original study, Thawani et al., (2020) benchmarked GP models against several alternatives on the Photoswitch dataset, including Random Forests, Attentive Neural Processes, Graph Convolutional Networks, Graph Attention Networks, and Directed Message Passing Neural Networks. While such models primarily target improvements in predictive accuracy within a standard regression framework, the focus of the present work is different: we aim to develop uncertainty-based sequential design strategies that efficiently guide data acquisition. Within this setting, GPs remain the natural modeling choice, as they provide closed-form analytical expressions for our acquisition criteria. This focus is not restrictive, as the best-performing model in the benchmark of Thawani et al., (2020) was a GP, and GPs consistently demonstrated strong performance in molecular property prediction (Deringer et al.,, 2021; Griffiths et al.,, 2023).
Within the GP framework, the choice of kernel heavily depends on the molecule representation used to encode chemical structures. We consider molecules represented by molecular fingerprints (Christie et al.,, 1993; Johnson and Maggiora,, 1990), which encode the presence or absence of characteristic substructures using binary digits. For fingerprint representations, the Tanimoto kernel (Gower,, 1971) is widely utilized (e.g., Ralaivola et al.,, 2005; Miyao and Funatsu,, 2019; Tripp et al.,, 2023), favored due to its simplicity, efficiency, and effectiveness in measuring the overlap between the presence or absence of molecular features. In addition to Morgan fingerprints, Thawani et al., (2020) explored alternative molecular representations, and determined that the best-performing combination was the GP model with a Tanimoto kernel and their own hybrid molecular descriptor called fragprints. However, in this work, we continue to use Morgan fingerprints, as they are more broadly known, while exploring a few alternative kernel choices in the supplementary material.
By combining the Tanimoto kernel with our targeted CRPS-based criteria, we demonstrate how scoring rules can be leveraged for efficient sequential design towards uncovering molecules of interest. The novel closed-form expressions for CRPS-based acquisition functions presented in this paper are broadly applicable beyond chemistry and open new avenues for advancing sequential design methodologies based on scoring rules. The structure of the remaining paper is as follows: Section 2 reviews the fundamentals of GP modeling, excursion set estimation and targeted sequential design. While Section 3 is dedicated to the threshold-weighted CRPS and the development of the acquisition criteria based thereon, the numerical tests are introduced in Section 4 and the results are presented in Section 5. Section 6 wraps up the main body with a discussion and concluding words.
@@ 2.1 Gaussian process modeling and excursion set estimation
In GP modeling, an unknown function $f:\mathcal{X}\rightarrow\mathbb{R}$ is assumed to be a realization of a GP $\boldsymbol{\xi}$. Here, we assume we observe noisy evaluations of $f$ at $n$ input locations $\{\boldsymbol{x}_{1},\dots,\boldsymbol{x}_{n}\}\subset\mathcal{X}$, yielding the observations: $Z_{i}=f(\boldsymbol{x_{i}})+\varepsilon_{i},\ \text{where }\varepsilon_{i}\sim\mathcal{N}(0,\tau^{2})\ \text{i.i.d.}$. Conditioning the GP on the observation event denoted as $\mathcal{A}_{n}=\{\boldsymbol{Z}_{n}=\boldsymbol{z}_{n}\}$ results in a posterior Gaussian distribution characterized by its predictive mean $m_{n}(\boldsymbol{x})$ and posterior covariance function $k_{n}(\boldsymbol{x},\boldsymbol{x^{\prime}})$, whose expressions are recalled in the supplementary material (Part A).
@@ 3.1 Threshold-weighted CRPS
Weighted scoring rules enable a targeted assessment of forecasts. The threshold-weighted CRPS (Matheson and Winkler,, 1976; Gneiting and Ranjan,, 2011) extends the CRPS by considering a weighting measure $\gamma$, which emphasizes the values of interest:
$\mathrm{CRPS}_{\gamma}(F,y)=\int_{-\infty}^{\infty}[F(u)-\mathbbm{1}\{y\leq u\}]^{2}\,\mathrm{d}\gamma(u).$ (7)
While $\gamma$ may potentially be taken as any Borel measure on $\mathbb{R}$, in our excursion setting, we consider two types of weighting measures (characterized by their densities with respect to the Lebesgue measure $\lambda$). The first weighting measure $\gamma_{1}$ only considers values above the threshold $t$ and is characterized by $\mathrm{d}\gamma_{1}(u)=\mathbbm{1}\{t\leq u\}\mathrm{d}\lambda(u)$. The second weighting measure $\gamma_{2}$ admits a Gaussian probability density centered on the threshold $t$ and with variance $\sigma_{\gamma}^{2}>0$: $\mathrm{d}\gamma_{2}(u)=\frac{1}{\sigma_{\gamma}}\phi\left(\frac{u-t}{\sigma_{\gamma}}\right)\mathrm{d}\lambda(u)$, where $\phi:\mathbb{R}\rightarrow(0,\infty),\phi(v):=\frac{1}{\sqrt{2\pi}}\exp^{-\frac{v^{2}}{2}}$.
@@
For a predictive distribution $\Phi_{\mu,\sigma^{2}}$, it holds with weighting measure $wI$ that
$\mathrm{CRPS}_{\gamma_{1}}(\Phi_{\mu,\sigma^{2}},y)=\sigma\left[-\widetilde{t}\Phi(\widetilde{t})^{2}+\widetilde{y^{t}}\left(2\Phi(\widetilde{y^{t}})-1\right)+\left(2\phi\left(\widetilde{y^{t}}\right)-2\phi\left(\widetilde{t}\right)\Phi(\widetilde{t})\right)-\frac{1}{\sqrt{\pi}}\left(1-\Phi(\widetilde{t}\sqrt{2})\right)\right],$
@@
For the weighting measure $\gamma_{2}$ (with mean $t$ and variance $\sigma_{\gamma}^{2}$), one obtains
$\mathrm{CRPS}_{\gamma_{2}}(\Phi_{\mu,\sigma^{2}},y)$ $=\Phi\left(\begin{pmatrix}0\\ 0\end{pmatrix};\begin{pmatrix}\mu-t\\ \mu-t\end{pmatrix},\begin{pmatrix}\sigma_{\gamma}^{2}+\sigma^{2}&\sigma_{\gamma}^{2}\\ \sigma_{\gamma}^{2}&\sigma_{\gamma}^{2}+\sigma^{2}\end{pmatrix}\right)$
$\quad\quad-2\Phi\left(\begin{pmatrix}0\\ 0\end{pmatrix};\begin{pmatrix}\mu-t\\ y-t\end{pmatrix},\begin{pmatrix}\sigma_{\gamma}^{2}+\sigma^{2}&\sigma_{\gamma}^{2}\\ \sigma_{\gamma}^{2}&\sigma_{\gamma}^{2}\end{pmatrix}\right)+\Phi\left(\frac{t-y}{\sigma_{\gamma}}\right),$
@@
See supplementary material (Part C.1). ∎
For illustration purposes, Figure 1 depicts the discussed versions of the CRPS for a Gaussian $\Phi_{\mu,\sigma^{2}}$ (solid line), showing the area under the integral for the CRPS (a and d), $\mathrm{CRPS}_{\gamma_{1}}$ (b and e), and $\mathrm{CRPS}_{\gamma_{2}}$ (c and f) across two scenarios: In the first row (a-c) we consider $y<t<\mu$, referring to a point which does not lie in the excursion set ($y<t$), but is falsely predicted to be in the excursion set ($\mu>t$, see Eq. 2). In the second row (d-f), we use $\mu<t<y$, such that the point does lie in the excursion set ($y>t$), but is falsely predicted to be not ($\mu<t$). Notably, both the (unweighted) CRPS and $\mathrm{CRPS}_{\gamma_{2}}$ are symmetric, yielding the same value for both scenarios. Contrarily, the $\mathrm{CRPS}_{\gamma_{1}}$ has a higher value in the second scenario.
Figure 1: Illustrations of the $\mathrm{CRPS}\left(\Phi_{\mu,\sigma^{2}},y\right)$ and $\mathrm{CRPS}_{\gamma}\left(\Phi_{\mu,\sigma^{2}},y\right)$ for a Gaussian CDF $\Phi_{\mu,\sigma^{2}}(u)$ (solid line), an observation $y$ (indicator line) and a threshold $t$. In the first row (a-c) we have $y=-3,t=0,\mu=3$ and in the second row (d-f) we have $\mu=-3,t=0,y=3$; we use $\sigma=2$ for the predictive Gaussian. The individual plots display the area under the integral for the CRPS (a and d), $\mathrm{CRPS}_{\gamma_{1}}$ (b and e) and $\mathrm{CRPS}_{\gamma_{2}}$ (c and f; $\sigma_{\gamma}=1.5$), with the dashed lines indicating the weighting measures $\gamma_{1}$ and $\gamma_{2}$. The numbers refer to the value and expected value of the respective criterion.
@@ 3.2 Pointwise selection criteria
Given that the threshold-weighted CRPS evaluates both accuracy and uncertainty of the GP model with respect to the excursion set of interest, it is a promising basis for an acquisition criterion. For a pointwise criterion, we consider the threshold-weighted CRPS of all the candidate points, and add the one to the training set which has the highest score (Eq. 3). However, in its original form for evaluation (Eq. 7), the threshold-weighted CRPS relies on the materialized value $y$ in the candidate point, which is unavailable during data acquisition. Therefore, we consider an expected version, where the expectation is taken with respect to the current predictive distribution. This is also known as the expected score or entropy function for proper scoring rules (Gneiting and Raftery,, 2007). The expected threshold-weighted CRPS with weighting measure $\gamma$ for predictive CDF $F$ is defined as,
$\mathrm{CRPS}_{\gamma}(F)=\int_{-\infty}^{\infty}\mathrm{CRPS}_{\gamma}(F,y^{\prime})\mathrm{d}F(y^{\prime}).$ (8)
The following theorem provides an alternative representation of this expected threshold-weighted CRPS and derives explicit formulas for Gaussian predictive distributions with weighting measures $\gamma_{1}$ and $\gamma_{2}$.
@@
For a predictive CDF $F(\cdot)$ and a weighting measure $\gamma$, it holds,
$\mathrm{CRPS}_{\gamma}(F)=\int F(u)\left(1-F(u)\right)\mathrm{d}\gamma(u).$
Particularly, for $F=\Phi_{\mu,\sigma^{2}}$ and $\mathrm{d}\gamma_{1}(u)=\mathbbm{1}\{t\leq u\}\mathrm{d}\lambda(u)$, one obtains,
[…]
@@
$\mathrm{CRPS}_{\gamma_{1}}(\Phi_{\mu,\sigma^{2}})=\sigma\mathbb{E}\left[\left(\widetilde{N}-\max\left(N,\frac{t-\mu}{\sigma}\right)\right)_{+}\right],$
for $N,\widetilde{N}\overset{i.i.d.}{\sim}\mathcal{N}(0,1)$. For $\mathrm{d}\gamma_{2}(u)=\frac{1}{\sigma_{\gamma}}\phi\left(\frac{u-t}{\sigma_{\gamma}}\right)\mathrm{d}\lambda(u)$, we get,
$\mathrm{CRPS}_{\gamma_{2}}(\Phi_{\mu,\sigma^{2}})$ $=\Phi\left(\begin{pmatrix}0\\ 0\end{pmatrix};\begin{pmatrix}\mu-t\\ t-\mu\end{pmatrix},\begin{pmatrix}\sigma_{\gamma}^{2}+\sigma^{2}&-\sigma_{\gamma}^{2}\\ -\sigma_{\gamma}^{2}&\sigma_{\gamma}^{2}+\sigma^{2}\end{pmatrix}\right)$.
@@
Supplementary material (Part C.2); the first result generalizes the corresponding expression for the traditional CRPS (Gneiting and Raftery,, 2007). ∎
To use this for pointwise selection criteria (Eq. 3), we can set $G_{n}(\boldsymbol{x})=\mathrm{CRPS}_{\gamma}(\Phi_{\mu,\sigma^{2}})$, where $\mu=m_{n}(\boldsymbol{x})$ and $\sigma^{2}=k_{n}(\boldsymbol{x,\boldsymbol{x}})$. To better understand the resulting behavior, we return to Figure 1, in which we also denote the expected score values. Both CRPS and $\mathrm{CRPS}_{\gamma_{2}}$ maintain symmetry in their expected scores (a, c, d, f), while $\mathrm{CRPS}_{\gamma_{1}}$ remains asymmetric but with a reversed orientation. In the original scoring rule relying on $y$, a missed point within the excursion set results in a higher score. In the expected score, a point mistakenly classified as being within the excursion set yields a higher expected score. In general, the $\mathrm{CRPS}_{\gamma_{1}}$ is higher for cases where $\Phi_{\mu,\sigma^{2}}$ has $m>t$, regardless of whether the prediction is correct (since the correct outcome is unknown). Figure 2a shows this relationship for $\Phi_{\mu,\sigma^{2}}$ with a fixed $t=0$. Both the values of $\mathrm{CRPS}_{\gamma_{1}}$ and $\mathrm{CRPS}_{\gamma_{2}}$ increase with higher uncertainty ($\sigma$) for the same $\mu$. However, while $\mathrm{CRPS}_{\gamma_{1}}$ continuously increases with rising $\mu$ when $\sigma$ is fixed, $\mathrm{CRPS}_{\gamma_{2}}$ remains symmetric around the threshold $t$ for varying $\mu$.
Figure 2: Illustrations of the expected $\mathrm{CRPS}_{\gamma_{1}}$ and $\mathrm{CRPS}_{\gamma_{2}}$ for a Gaussian $\Phi_{\mu,\sigma^{2}}$ depending on the predictive mean $\mu$: (a) Expected $\mathrm{CRPS}_{\gamma_{1}}$ and $\mathrm{CRPS}_{\gamma_{2}}$ ($\sigma_{\gamma}=1.5$) and (b) Expected $\mathrm{CRPS}_{\gamma_{2}}$ for different values of $\sigma_{\gamma}$ ($\sigma=2$). In both plots, the vertical solid line indicate the threshold $t=0$.
While the indicator density of $\mathrm{CRPS}_{\gamma_{1}}$ is fixed (although a shift would be thinkable), the density of $\mathrm{CRPS}_{\gamma_{2}}$ is influenced by the parameter $\sigma_{\gamma}$, which controls the width of the weight around the threshold $t$. Figure 2b illustrates the dependence of $\mathrm{CRPS}_{\gamma_{2}}$ on the predictive mean $\mu$ for different values of $\sigma_{\gamma}$ with $t=0$. A lower value of $\sigma_{\gamma}$ places greater emphasis on the region near the threshold, while a higher value results in more uniform values across different $\mu$.